# Quantumness and Randomness of Quantum Random Number Generators via Photon Statistics

### v2

@@ IV.4 Time-bin Statistics for Poissonian Photon with Erroneous Devices
We have performed the NIST test Bassham et al. (2010) on the data generated by simulating the probability distribution given by Equation 15. Test results are provided in Appendix B.
V Conclusion
@@
In this article, we assume that photon detection time is available as a sample. We discuss how to determine when the source can be considered as quantum. We suggest two-fold statistical tests. The first one may be an interval estimation or hypothesis test to decide between sub-Poissonian and super-Poissonian. On the other hand, the last one, which is a $\chi^{2}$-test ensures Poisson distribution. This test can be performed for any quantum-photonic device to check the quantumness in it. Then we consider photon statistics for time-bin-based QRNG in practical scenarios. We consider two photon arrival time-based QRNG models, one using external time reference and another without such external reference. We show that, although for ideal devices these two models follow different distributions (one of which is perfectly random), these two QRNGs follow a similar distribution when the photon source is not an actual single photon source. We derive a relation when the photon statistics becomes $\epsilon$-random. We found that by making the product of the expected number of photons ($\mu T$) within an external reference cycle ($T$) and the detector efficiency ($d$) small enough, the randomness of the generated raw random number can be increased. Therefore it would require less post-processing, and it would retain maximum randomness from the quantum source.
@@
In practical cases, the dark count of a detector also contributes to the generation of random numbers. In our discussion, we did not consider the dark count. This work can be extended to see the effect of dark counts on the photon statistics. One may note that our discussions are not meant for black-box testing for commercially available QRNGs. Rather, the physicists and engineers may use these in the QRNG design phase before post-processing, for evaluation, fine-tuning and performance optimization.
VI Acknowledgement
@@
where $\langle N\rangle$ and $\langle(\Delta N)^{2}\rangle$ are respectively the mean and the variance of the above distribution. Choosing $T=\frac{2\pi}{\omega_{1}-\omega_{2}}$, one can see that $\langle(\Delta N)^{2}\rangle<\langle N\rangle$, i.e., the light is sub-Poissonian although $g^{2}(0)=\frac{3}{2}-\frac{1}{n}>1$ for $n>2$ denoting the light is bunched. This example also shows that although both the antibunching of photons and the sub-Poissonian distribution of photon count indicate the non-classicality of light, they are not identical.
Appendix B Statistical Test of Randomness
As mentioned in Subsection IV.1, the randomness of the random numbers is tested using statistical tests. Different statistical testing suites like NIST Bassham et al. (2010), Dieharder Brown et al. (2014), ENT Walker (2008) etc. are available to serve this purpose. These suites consider different aspects, such as entropy, serial correlation, frequency of values, repetition of similar patterns etc. to perform the test. Some tests also perform Monte Carlo simulation Metropolis and Ulam (1949) to validate the randomness.
The NIST Randomness Testing Suite is an extensive set of statistical tests developed by the National Institute of Standards and Technology (NIST) to evaluate the quality of randomness in random numbers. It is one of the most widely used suites for testing randomness. It contains $15$ different tests to examine different aspects of randomness. Most of the tests have the standard normal or $\chi^{2}$ as reference distribution for their test statistic. Here null hypothesis is taken as the number to be tested is random. For each test, the significance level is considered as $0.01$. If the value of the test statistic for some tests falls in the critical region, or equivalently, calculated $p$-value is less than $0.01$, then the corresponding tests fail rejecting the null hypothesis. A perfectly random number should pass all of the $15$ tests. As mentioned in Section III, the central limit theorem is applicable only for large $n$, tests that use the central limit theorem to determine the reference distribution, have a requirement for the bit stream, that is, the length of the testing sequence, to be large enough (at least $10^{6}$).
Figure 7: NIST result. Bar charts are plotted for (a) $\mu Td=0.05$, (b) $\mu Td=0.10$, (c) $\mu Td=0.15$, (d) $\mu Td=0.20$, (e) $\mu Td=0.25$, (f) $\mu Td=0.30$, (g) $\mu Td=0.35$, (h) $\mu Td=0.40$. For each $\mu Td$, three $1$ GB files are generated by simulating probability distribution (15), and the NIST test has been performed on those files. Different textures denote results corresponding to different files. A test is considered as passed if the corresponding $p$-value is at least $0.01$. Tests that fail for all of the three files are not included in the bar chart. The vertical dashed line denotes $p$-value $=0.01$
Value of $\bm{\mu Td}$ Number of NIST tests passed
$0.05$ $10$-$12/15$
$0.10$ $8$-$10/15$
$0.15$ $8/15$
$0.20$ $6$-$7/15$
$0.25$ $6$-$7/15$
$0.30$ $5$-$6/15$
$0.35$ $6/15$
$0.40$ $4$-$5/15$
Table 1: NIST test result for the data simulated with the distribution given by Equation (15). Three files, each containing $1$ GB random number have been used for the test for every $\mu Td$. A test fails if the corresponding $p$-value is found less than $0.01$. The range in the second column arises due to different results from three different files. The number of passed tests decreases with the increment of $\mu Td$. This indicates that as $\mu Td$ increases, the randomness in the generated random number decreases.
We have generated random numbers simulating the probability distribution given by Equation (15) for different $\mu Td$. Three files of size $1$ GB have been generated for every $\mu Td$. Then we perform the NIST tests over the generated numbers. $10$ bit streams of length $10,000,000$ have been considered for the tests. Bar charts of $p$-values from different NIST tests are plotted in Figure 7. The tests for which the $p$-values are found as 00 are not included in the plot. From this plot, one can see that, as $\mu Td$ increases, the number of failed tests also increases indicating a decrement of the randomness in the random number. In Table 1, We tabulate the number of tests our simulated data passed. This table indicates that lower $\mu Td$ values give rise to better random numbers requiring less post-processing.

### v3

@@ VI.2 Time-bin Statistics for Poissonian Photon with Inperfect Devices
We have performed the NIST, Dieharder, AIS-31 and ENT tests on the data generated by simulating the probability distribution given by Eq. (29). These test results are provided in Section VIII.
VII Effect of $\mu,T,d$ on random number generation rate and cost
As discussed in Eq. (30) of Subsection VI.2, we have to minimize $\mu Td$ to achieve maximum randomness from the quantum source increases. In that discussion, we have considered the product $\mu Td$ as a single quantity. In this section, we separately discuss the effect of
  • the expected photon number in unit time, $\mu$,
  • the external reference cycle, $T$, and
  • the detection efficiency, $d$
on the rate of random number generation, $r$ as well as on the cost of the QRNG.
The entropy, which is the source of randomness of a QRNG, can be expressed as
$H(E)=H(Q)+H(C),$ (35)
where, $H(E),H(Q)$ and $H(C)$ are the total entropy generated by the RNG, the entropy of the quantum source and the entropy of the classical source respectively.
Definition 3.
An RNG is said to be perfect if $H(E)=H(U)$, where $H(U)$ is the entropy of uniform randomness. A QRNG is said to be perfect if $H(U)=H(E)=H(Q)$.
However, in the practical scenario, perfect QRNG is not possible. Therefore, the main goal in designing a QRNG is
[…]
Figure 7: Entropy (38) of the probability distribution (29). When the product of the expected number of photons, $\mu T$, within an external reference cycle $T$, and the detection efficiency $d$ is very small, the entropy $H(Q)$ of the distribution generated by the QRNG is very close to the entropy $H(U)=8$ of the uniform random distribution. Therefore, if $\mu Td$ is very small, the random number generated by the QRNG would be close to the uniform random number.
The Shannon entropy Shannon (1948a, b) of a uniform distribution over $N$ points is given by
$H(U)=-\sum_{i=1}^{N}\frac{1}{N}\log\frac{1}{N}=\log N.$ (37)
On the other hand, the entropy of the probability distribution (29) is given by
$H(Q)=\frac{1-(N+1)e^{-\mu Td}+Ne^{-(N+1)\frac{\mu Td}{N}}}{(1-e^{-\mu Td})(1-e^{-\frac{\mu Td}{N}})}\log\left(e^{\frac{\mu Td}{N}}\right)-\log\left(\frac{e^{\frac{\mu Td}{N}}-1}{1-e^{-\mu Td}}\right).$ (38)
A detailed derivation of this entropy has been added in Appendix B. Fig. 7 shows how the entropy (38) changes with $\mu Td$. Here we have chosen $N=256$. Therefore, the entropy (37) of the uniform distribution would be $H(U)=\log 256=8$. From Fig. 7, it can be easily seen that, if $\mu Td$ is very small, the entropy $H(Q)$ is very close to $8$, which is the entropy of the uniform distribution. In this section, by randomness we will refer to the randomness due to $H(Q)$ only.
VII.1 Effect of the Expected Photon Numbers
Here we discuss the effect of the expected photon numbers on the cost as well as the random number generation rate of a QRNG.
VII.1.1 Effect of the Exected Photon Number on Cost
Since $\mu$ is the expected number of photons in unit time, this can be reduced by using heavy attenuation. This will directly affect the cost. Heavy attenuation requires a large number of high-quality attenuators, increasing the production cost for the random numbers. Thus
small value of μ high production $\text{small value of $\mu\implies$ high production cost}.$ (39)
VII.1.2 Effect of the Expected Photon Number on Random Number Generation Rate
A small value of $\mu$, which indicates a small number of photons, leads to a high chance of no photon detection in a particular external reference cycle of time interval $T$. That means those intervals will have no contribution to the final random number. Therefore, we require more cycles to produce random numbers of expected length. This will reduce the speed of the RNG. From Eq. (VI.2), the probability, $P(0)$, of no photon detection within time $T$ is given by
$P(0)=1-\sum_{i=1}^{N}P_{d}(i;1st)=e^{-\mu Td},$ (40)
where $P_{d}(i;1st)$ is the probability of detecting the first photon in the $i$-th bin.
Figure 8: Plot shows how random number generation rate (bit per unit time) changes with the expected photon number per unit time ($\mu$). Five different plots have been shown for five different values of detection efficiency ($d$). From the plot, it can be easily seen that if $\mu$ decreases, the random number generation rate of the QRNG will also decrease.
Suppose, for some $\mu$, the expected number of external cycles to produce a random number of length $k$, where each cycle contains $N$ time-bins, is given by $N_{c}$. Then $N_{c}=\langle X\rangle$, where $X$ is a random variable, whose probability distribution is negative binomial distribution, given by
$P(X=x)=\binom{x-1}{k^{\prime}-1}\left(1-P_{\mu}(0)\right)^{k^{\prime}}P_{\mu}(0)^{x-k^{\prime}},$ (41)
for $x=k^{\prime},k^{\prime}+1,k^{\prime}+2,\dots$, where $k^{\prime}=\lceil\frac{k}{\log_{2}N}\rceil$ and $P_{\mu}(0)=e^{-\mu Td}$. Therefore,
$N_{c}=\langle X\rangle=\frac{k^{\prime}P(0)}{1-P(0)}=\frac{k^{\prime}}{e^{\mu Td}-1}.$ (42)
Thus, the rate of generating random numbers is given by
$r=\frac{k^{\prime}}{TN_{c}}=\frac{e^{\mu Td}-1}{T}\text{ bits per unit time}.$ (43)
Since the external time cycle $T$ is very small, Eq. (43) can be written as
$r\approx\mu d\text{ bits per unit time}.$ (44)
Clearly, if $\mu$ decreases, this rate will also decrease. The graph in Fig. 8 shows how this rate changes with $\mu$ for fixed $T$ and $d$. Therefore, we cannot reduce $\mu$ arbitrarily to get the maximum amount of randomness.
VII.2 Effect of the timing error of the device on the External Reference Cycle
If we choose the external reference cycle $T$ very small, the product $\mu Td$ would also be small. This should increase the randomness, as discussed in Eq. (30) of the Subsection VI.2. However, a small $T$ would lead to a small length $\frac{T}{N}$ for the time bins. Now, if there are maximum $\delta_{t}$ errors in registering the photon-arrival-time, due to small time-bins, the time may be registered in a time-bin that is different from the actual time-bin as shown in Fig. 9. Let us consider that the actual detection happens in $i$-th bin, that is, within time interval $[(i-1)\frac{T}{N},i\frac{T}{N})$. If the actual detection happens within time interval $[(i-1)\frac{T}{N},(i-1)\frac{T}{N}+\delta_{t})$, the time may be registered as $(i-1)$-th bin. Similarly, if the actual detection happens within time interval $[i\frac{T}{N}-\delta_{t},i\frac{T}{N})$, the time may be registered as $(i+1)$-th bin. However, if the actual detection lies within the time interval $[(i-1)\frac{T}{N}+\delta_{t},i\frac{T}{N}-\delta_{t})$, there will be no error.
[…]
Figure 9: Due to the timing error $\delta_{t}$, the actual may be read as the shifted signal and the registered time-bin may be different from the actual time-bin of photon detection. This would add an error in the generated random number. To minimize such error, we put a lower bound on the external reference cycle $T$ using Eq. (46).
Now let us consider that $T$ is chosen to satisfy the relation
$\frac{T}{N}\geq k\delta_{t},$ (45)
for some suitable $k$. Then the probability, $p_{e}$, that the detection happens in $i$-th bin but outside of the interval $[(i-1)\frac{T}{N}+\delta_{t},i\frac{T}{N}-\delta_{t})$ is bounded above as $p_{e}\leq\frac{2}{k}$. Here we have assumed uniform distribution as the time $\frac{T}{Nk}$ is very small. Therefore, for a fixed error tolerance, $p_{tol}$, $k$ can be chosen as $k_{tol}=\frac{2}{p_{tol}}$. In that case, the minimum reference cycle $T_{min}$ should be
$T_{min}=Nk_{tol}\delta_{t}.$ (46)
Therefore, we cannot make $T$ arbitrarily small to get the maximum amount of randomness.
VII.3 Effect of the Reference Cycle on Random Number Generation Rate and Cost
Equation (44) indicates that $T$ does not have much effect on the rate of generating random numbers. Also, since $T$ has no impact on the used optical devices, it does not affect cost.
VII.4 Effect of Detection Efficiency
In this subsection, we discuss the effect of the detection efficiency of the detector used in the QRNG on the cost as well as the random number generation rate of a QRNG.
VII.4.1 Effect of the Detection Efficiency on Cost
The value of $\mu Td$ can be decreased by choosing a detector with bad detection efficiency. It is very natural that if a detector can efficiently detect photons, its cost will also be high. Since the randomness increases if $\mu Td$ decreases, choosing small $d$ we can increase the randomness. This will also reduce the cost of the QRNG. Thus
small value of $\text{small value of }d\implies\begin{cases}\text{high randomness}\\ \text{high cost}\end{cases}$ (47)
VII.4.2 Effect of the Detection Efficiency on Random Number Generation Rate
Figure 10: Plot shows how random number generation rate (bit per unit time) changes with the detection efficiency ($d$). Five different plots have been shown for five different values of $\mu$, the expected photon number per unit time. From the plot, it can be easily seen that if $d$ decreases, the random number generation rate of the QRNG will also decrease.
Although a bad detector with a small detection efficiency, $d$, reduces production costs, this leads to a high chance of no photon detection in a particular external reference cycle. From Eq. (44) it can be easily seen that if $T$ decreases, the rate of generating random numbers will also decrease. The graph in Fig. 10 shows how this rate changes with $\mu$ for fixed $T$ and $d$. Therefore, it is not possible to choose a detector with arbitrarily small detection efficiency.
VII.5 Performance and Cost of the QRNG
Figure 11: Plot showing relation between random number generation rate, $r$, expected photon number in unit time, $\mu$, and detection efficiency, $d$. If the values of $\mu$ and $d$ are small, the generation rate is also small. Therefore, to generate random numbers with high rate, we have to choose both $\mu$ and $d$ large.
Figure 12: Heat map denoting how (a) randomness, (b) random number generation rate and (c) cost varies depending on the expected photon number per unit time, $\mu$, external reference cycle, $T$ and the detection efficiency, $d$. Figure (a) shows that we must keep the product $\mu Td$ small to get maximum randomness. However, (b) shows if the $\mu d$ is small, the generation rate will be minimal. On the other hand, from (c), low-cost demands a small value of $d$ and a high value of $\mu$.
Figure 11 is a surface plot of how the random number generation rate of the QRNG changes with $\mu$ and $d$. This shows that we have to consider high values of $\mu$ and $d$ to get a high rate. On the other hand, in Eq. (30) of Subsection VI.2 we show that to get maximum randomness from the quantum source, we have to minimize the product $\mu Td$ small. Therefore, this needs some optimization while choosing $\mu$ and $d$.
Parameters Effects on
$\downarrow$ Randomness Rate Cost
$\mu$uparrow $\downarrow$ $\uparrow$
$T$uparrow -- --
$d$uparrow $\downarrow$ $\downarrow$
Table 1: The effects of decreasing the expected photon count $\mu$, the external reference cycle $T$ and the detection efficiency $d$. A down-arrow, $\downarrow$, denotes the decrement of the corresponding quantity. On the other hand, an up-arrow, $\uparrow$, denotes the increment. $T$ does not affect speed or cost. If the parameters increase, the effects will be in the reverse direction.
The above discussion has been described as a heat map in Fig. 12. In the figure, dark shade denotes low value and light shade denotes high value of randomness, generation rate and cost in Figures 12(a), 12(b) and 12(c) respectively. Figure 12(a) shows that high randomness demands a small value of the product $\mu Td$. However, the product $\mu d$ cannot be very small as Fig. 12(b) indicates. A small $\mu d$ would result in a low random number generation rate. On the other hand, to keep the cost of the QRNG low, we need a small $d$ but a large $\mu$.
In Table 1, we summarize the discussion in tabular form. A down-arrow ($\downarrow$) indicates a decrease in the value of the corresponding quantity, while an up-arrow ($\uparrow$) signifies an increase. The external reference cycle $T$ does not influence speed or cost. If the parameters are increased, the effects will reverse accordingly.
We have already mentioned in Eq. (36) that the main goal in designing a QRNG is to keep $H(Q)$, the entropy from the quantum source as close as $H(U)$, the entropy of uniform randomness. However, we also discussed that this is not the only goal. There are some other constraints as well like random number generation rate and cost. A QRNG should generate random numbers with high rates and low costs. The randomness and the rate together describe the performance of the QRNG. Therefore, the goal in designing a QRNG is to maximize the performance and minimize the cost by choosing appropriate $\mu,T$ and $d$.
VIII Results of Randomness Tests
As mentioned in Subsection II.3, the randomness of the random numbers is tested using statistical tests. Here we used four common statistical testing suites namely NIST, Dieharder, AIS-31 and ENT.
Figure 13: NIST result. Bar charts are plotted for (a) $\mu Td=0.05$, (b) $\mu Td=0.10$, (c) $\mu Td=0.15$, (d) $\mu Td=0.20$, (e) $\mu Td=0.25$, (f) $\mu Td=0.30$, (g) $\mu Td=0.35$, (h) $\mu Td=0.40$, (i) $\mu Td=0.45$, (j) $\mu Td=0.50$. For each $\mu Td$, three $1$ GB files are generated by simulating probability distribution (29), and the NIST tests have been performed on those files. Different textures denote results corresponding to different files. A test is considered as passed if the corresponding $p$-value is at least $0.01$. Tests that fail for all of the three files are not included in the bar chart. The vertical dashed line denotes $p$-value $=0.01$
Figure 14: NIST test results for the data simulated with the distribution given by Eq. (29). Three files, each containing $1$ GB random number have been used for the test for every $\mu Td$. A test fails if the corresponding $p$-value is found less than $0.01$. The range in the second column arises due to different results from three different files. The number of passed tests decreases with the increment of $\mu Td$. This indicates that as $\mu Td$ increases, the randomness in the generated random number decreases.
The NIST Randomness Testing Suite is an extensive set of statistical tests developed by the National Institute of Standards and Technology (NIST) to evaluate the quality of randomness in random numbers. It is one of the most widely used suites for testing randomness. It contains $15$ different tests to examine different aspects of randomness. Most of the tests have the standard normal or $\chi^{2}$ as reference distribution for their test statistic. Here null hypothesis is taken as the number to be tested is random. For each test, the significance level is considered as $0.01$. If the value of the test statistic for some tests falls in the critical region, or equivalently, calculated $p$-value is less than $0.01$, then the corresponding tests fail rejecting the null hypothesis. A perfectly random number should pass all of the $15$ tests. As mentioned in Section V, the central limit theorem is applicable only for large $n$, tests that use the central limit theorem to determine the reference distribution, have a requirement for the bit stream, that is, the length of the testing sequence, to be large enough (at least $10^{6}$).
The Dieharder tests represent a battery of statistical tests used to assess the quality and randomness of a RNG. This suite contains around $30$ tests for the assessment. These tests consider KS test of uniformity to obtain the $p$-values. Depending on these $p$-values, the decision (passed, weak or failed) has been taken. The results that do not pass and also do not have enough evidence to declare as fail have been mentioned as weak.
AIS-$31$ is a standardized suite of statistical tests for assessing RNGs. This suite comprises a total of 9 tests, typically organized into two main groups. Procedure A, consists of 6 tests aimed at detecting statistically inconspicuous behavior in RNG output. Procedure B, focuses on evaluating the internal random bits of TRNGs through 3 specific tests. These tests ensure that RNGs meet rigorous statistical criteria for randomness, providing a reliable measure of their quality and suitability for applications requiring secure and unpredictable random numbers.
ENT performs $5$ tests on the stream of bytes and produces the outputs for each test. It calculates entropy, $\chi^{2}$ statistic value, the arithmetic mean of the bytes, the Monte Carlo value of $\pi$ and the serial correlation coefficient. Depending on these values user can decide the randomness.
Figure 15: (a) Dieharder and (b) AIS-$31$ test results for the data simulated with the distribution given by Eq. (29). For Dieharder, there are $30$ tests; for AIS-$31$, $9$ tests are there. These tests also reveal that the number of passed tests decreases as $\mu Td$ increases indicating lower values of $\mu Td$ result in better randomness before post-processing. Along with pass and fail, Dieharder tests also mention suspected results as weak. They are also included in this figure.
We have generated random numbers simulating the probability distribution given by Eq. (29) for different $\mu Td$. Three files of size $1$ GB have been generated for every $\mu Td$. Then we perform the NIST tests over the generated numbers. $100$ bit streams of length $10,000,000$ have been considered for the tests. In Fig. 13, we plot the bar charts of the $p$-values from different NIST tests. The tests for which the $p$-values are found as 00 are not included in the plot. From these plots, we can see that, as $\mu Td$ increases, the number of passed tests decreases indicating a decrement of the randomness in the random number. The overall result is plotted in Fig. 14 as the number of tests our simulated data passed. These graphs indicate that lower $\mu Td$ values give rise to better random numbers requiring less post-processing.
Figure 16: Ent test result for the data simulated with the distribution given by Eq. (29). The relative deviation is calculated as $\frac{|v_{i}-v_{o}|}{v_{i}}$, where $v_{i}$ is the expected ideal value for uniform distribution and $v_{o}$ is the observed value for the simulated data. Although the deviation of the Monte Carlo value of $\pi$, the arithmetic mean and the $\chi^{2}$ test value is high, the deviation of the entropy and the serial correlation is very small.
We have also performed the Dieharder, AIS-$31$ and ENT over the generated numbers. Figure 15 shows how many tests for our simulated data passed for Dieharder and AIS tests. Similar to the NIST tests, these tests also show that small $\mu Td$ values provide better randomness before post-processing. ENT test does not mention pass or fail as a result. However, it provides ideal and actual values for the performed tests. In Fig. 16, we plot the relative deviation of the actual values from their ideal values for different ENT tests. In this case also, the deviations increase with the increase of $\mu Td$.
IX Discussion
@@
In this article, we consider two perspectives for analyzing QRNG. From the manufacturer’s point of view, we discuss how to determine when the source can be considered as quantum. Here we assume that photon detection time is available as a sample. We suggest two-fold statistical tests. The first one may be an interval estimation or hypothesis test to decide between sub-Poissonian and super-Poissonian. On the other hand, the last one, which is a $\chi^{2}$-test ensures Poisson distribution. This test can be performed for any quantum-photonic device to check its quantumness.
Then we discuss photon statistics for time-bin-based QRNG in practical scenarios. We consider two photon-arrival-time-based QRNG models, one using external time reference and another without such external reference. We show that, although for ideal devices these two models follow different distributions (one of which is perfectly random), these two QRNGs follow a similar distribution when the photon source is not an actual single photon source. We derive a relation when the photon statistics becomes $\epsilon$-random. We found that by making the product of the expected number of photons ($\mu T$) within an external reference cycle ($T$) and the detection efficiency ($d$) small enough, the randomness of the generated raw random number can be increased. Therefore it would require less post-processing, and it would retain maximum randomness from the quantum source. This discussion would help QRNG manufacturers to design better QRNG.
Although small $\mu Td$ provides better randomness, from the user’s perspective it may not be suitable. The random number generation rate, the timing error in registering photon detection time and the cost of the QRNG restrict us from choosing small values for $\mu,T$ or $d$. A lower expected photon count improves quantumness but comes with increased costs and a reduced generation rate. On the other hand, a shorter external reference cycle enhances quantumness but must surpass a minimum threshold to reduce timing errors, having only slight effects on cost and rate. Reduced detection efficiency boosts quantumness and lowers costs, but also decreases the generation rate. The trade-off among the randomness sourced from quantum systems, the rate of random number generation, and the cost of the QRNG is contingent upon the specific requirements of the users.
Finally, we simulated random numbers according to the photon statistics we derived and applied statistical tests including NIST, Dieharder, AIS-31, and ENT to these numbers. The results from these tests validate our theoretical conclusions based on $\mu Td$ about the randomness of the quantum sources.
In practical cases, the dark count of a detector also contributes to the generation of random numbers. In our discussion, we did not consider the dark count. This work can be extended to see the effect of dark counts on the photon statistics. In our discussion, we have considered two models of photon arrival-time-based QRNGs. Similar analysis can be performed for other models of QRNGs.
@@
One may note that our discussions are not meant for black-box testing for commercially available QRNGs. Rather, the physicists and engineers may use these in the QRNG design phase before post-processing, for evaluation, fine-tuning and performance optimization.
Acknowledgement
@@
where $\langle N\rangle$ and $\langle(\Delta N)^{2}\rangle$ are respectively the mean and the variance of the above distribution. Choosing $T=\frac{2\pi}{\omega_{1}-\omega_{2}}$, it can be seen that $\langle(\Delta N)^{2}\rangle<\langle N\rangle$, i.e., the light is sub-Poissonian although $g^{2}(0)=\frac{3}{2}-\frac{1}{n}>1$ for $n>2$ denoting the light is bunched. This example also shows that although both the antibunching of photons and the sub-Poissonian distribution of photon count indicate the non-classicality of light, they are not identical.
Appendix B Quantum Entropy for QRNG
The Shannon entropy Shannon (1948a, b) of the probability distribution given by Eq. (29) is given by
$H(Q)$ $=-\sum_{i=1}^{N}f(i)\log f(i)$
$=-\sum_{i=1}^{N}\frac{e^{\frac{\mu Td}{N}}-1}{1-e^{-\mu Td}}e^{-i\frac{\mu Td}{N}}\log\left(\frac{e^{\frac{\mu Td}{N}}-1}{1-e^{-\mu Td}}e^{-i\frac{\mu Td}{N}}\right)$
$=-\frac{e^{\frac{\mu Td}{N}}-1}{1-e^{-\mu Td}}\sum_{i=1}^{N}e^{-i\frac{\mu Td}{N}}\left[\log\left(e^{-i\frac{\mu Td}{N}}\right)+\log\left(\frac{e^{\frac{\mu Td}{N}}-1}{1-e^{-\mu Td}}\right)\right]$
$=\frac{e^{\frac{\mu Td}{N}}-1}{1-e^{-\mu Td}}\left[\log\left(e^{\frac{\mu Td}{N}}\right)\sum_{i=1}^{N}ie^{-i\frac{\mu Td}{N}}-\log\left(\frac{e^{\frac{\mu Td}{N}}-1}{1-e^{-\mu Td}}\right)\sum_{i=1}^{N}e^{-i\frac{\mu Td}{N}}\right]$
$=\frac{e^{\frac{\mu Td}{N}}-1}{1-e^{-\mu Td}}\left[\log\left(e^{\frac{\mu Td}{N}}\right)\frac{e^{-\frac{\mu Td}{N}}\left[1-(N+1)e^{-\mu Td}+Ne^{-(N+1)\frac{\mu Td}{N}}\right]}{(1-e^{-\frac{\mu Td}{N}})^{2}}-\log\left(\frac{e^{\frac{\mu Td}{N}}-1}{1-e^{-\mu Td}}\right)\frac{e^{-\frac{\mu Td}{N}}(1-e^{-\mu Td})}{1-e^{-\frac{\mu Td}{N}}}\right]$
$=\frac{1-(N+1)e^{-\mu Td}+Ne^{-(N+1)\frac{\mu Td}{N}}}{(1-e^{-\mu Td})(1-e^{-\frac{\mu Td}{N}})}\log\left(e^{\frac{\mu Td}{N}}\right)-\log\left(\frac{e^{\frac{\mu Td}{N}}-1}{1-e^{-\mu Td}}\right).$ (54)
This is the entropy generated by the photon arrival-time-based QRNGs we have discussed in this article.